OpenStudy (anonymous):

By recognizing the series below as a Taylor series evaluated at a particular value of x, find the exact sum of the convergent series.

OpenStudy (anonymous):

\[1-\frac{ 81 }{ 2! }+\frac{ 6561 }{ 4! }+...+(-1)^n \frac{ 9^{2n} }{ (2n)! }+...\]